OpenStudy (anonymous):

how do you simplify √9 over 6√9?

Nnesha (nnesha):

okay what is square root of 9 ??

OpenStudy (anonymous):

3

OpenStudy (anonymous):

i got 3 over 18 but none of those are the multiple choice answers

Nnesha (nnesha):

\[\huge\rm \frac{ \color{green}{\sqrt{9}} }{ 6\color{green}{\sqrt{9}} }\] \[\frac{ \color{green}3 }{ 6 \times \color{green}3 }\] can't you cancel out 3 ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

so cancel out 3 and then you will get your answer

OpenStudy (anonymous):

but 6 isnt one of the choices

Nnesha (nnesha):

how did you get 6 ??

OpenStudy (anonymous):

its 9 -2/3, 9 -1/2, 9 1/2 and 9 1/3 are the possible answers

Nnesha (nnesha):

are u sure those are answer choices for that questions ?

OpenStudy (anonymous):

yes